OpenStudy (yash2651995):

what does RMS (root mean square) things signify? V rms (rms velocity)? I rms (current in ac)? V rms (voltage in ac)?

OpenStudy (ghazi):

RMS stands for root mean square...for any case from thermodynamics to electricity and every where it is defined as the square root of the mean of that quantity over a time interval....now first of all to find the mean of that quantity you need to integrate the square of the function of that quantity over a given time interval and then take the square root of it..you'll get rms ....like \[I rms = \sqrt{1/(T2-T1)* \int\limits_{T1}^{T2} (Ip \sin wt)^{2}} dt\] here Ip is the peak amplitude of current that has sinusoidal variation ...T2, T1 is the time interval over which you've to find rms..similarly you can calculate rms for other quantities

OpenStudy (yash2651995):

but what does it signify? what it shows..? just a quantity? just a way of measurement? or it signifies something?

OpenStudy (ghazi):

Electrical engineers often need to know the power, dissipated by an electrical resistance. It is easy to do the calculation when there is a constant current, , through the resistance. For a load of ohms, power is defined simply as: P=I^2*R However, if the current (or any time varying function) is a time-varying function,I(t) ,this formula must be extended to reflect the fact that the current (and thus the instantaneous power) is varying over time. If the function is periodic (such as household AC power), it is nonetheless still meaningful to talk about the average power dissipated over time, which we calculate by taking the simple average of the power at each instant in the waveform or, equivalently, the squared current. That is, P=R*(I(t)rms)^2 So, the RMS value,of the function is the constant signal that yields the same power dissipation as the time-averaged power dissipation of the current . So basically, RMS value is used for time varying function.

OpenStudy (yash2651995):

ok.. it a type of average value.. my teacher said that rms values are those values of current in AC which correspond to the power dissipated in the DC.. ie if, i rms of a given AC current =k amperes then, power by the AC current in the given circuit is the same amount of power in same amount of time is obtained (in same circuit, in which AC current is flowing) if DC current of magnitude k amperes is flowing in it. (i think, i might have messed up.. what he told was convincing.. i forgot what he exactly said.. it might be wrong here.. i should have noted it down)

OpenStudy (ghazi):

basically...we use this method when the function of voltage or current is varying with time...else no need to use rms :)